OpenStudy (anonymous):

Circle P has a radius of 8 units with center P at (2, -1). Which equation defines circle P?

OpenStudy (anonymous):

standard eq for a circle is (x - h)^2 + (y-k)^2 = r^2 where (h,k) = center of the circle r = radius. just plug in the values and u'll get the eq.

OpenStudy (anonymous):

64=(x-2)^2+(y+1)^2

OpenStudy (anonymous):

thanks :)

OpenStudy (anonymous):

 r^2=y^2+x^2 is from Pythagoras